On January 1, 2019, John Corp. sold a parcel of land to Goodman Inc. John had purchased the land three years ago at $110,000. John received $50,000 cash and a noninterest bearing note for $320,000 from Goodman Inc. The note is due December 31, 2021 There is no readily available market value for the land, but the current market rate for comparable notes is 8%. SHOW YOUR LABELLED WORK!

1-1. Determine the selling price of the land.

1-2. Record the sale of the land by John .

1-3. Prepare an amortization table for John using the effective interest rate method.

1-4. Prepare any necessary journal entry(ies) at the end of 2020.

1-5. How would the transactions relating to this situation be reported on Johns December 31, 2019 statement of cash flows?
Answer #1

1-1) Selling Price of land:

Cash:$50000

Interest on note for 2 years: $320000*8%*2=51200

Note Value: $320000

Therefore Selling price of land= 50000+320000+51200= $421200

1-2)Bank A/c Dr. $50000

Promissory note A/c Dr. $320000

To land A/c $110000

To profit on sale of land A/c $260000

(being asset sold and profit booked)

1-4) Profit on sale of land A/c $260000

To p/l $260000

(being profit transferred to pl/l a/c)

1-5) cash flow from Investing activities will increase by $50000

1-4)